OpenStudy (anonymous):

The sum of three consecutive integers is 120. What are the integers. Find three consecutive integers such that the sum of the first and third equals the sum of the second and 21 Find four consecutive even integers such that four times the sum of the first and second is 8 less tgab 7 times the fourth And i have no idea about any of the steps so...

OpenStudy (anonymous):

if x is the first integer, the next consecutive integer is x+1 and the one after that is x+1+1 their sum is 120 so x +x+1+x+1+1 = 120 3x = 117 x = 39

OpenStudy (anonymous):

-The sum of three consecutive integers is 120. What are the integers. Well, the average value of the three numbers must be 120/3 = 40 so, 39, 40, 41 -Find three consecutive integers such that the sum of the first and third equals the sum of the second and 21 so x + (x + 2) = (x + 1) + 21 Solving for x will give 20 as the first number, so 20, 21, and 22. -Find four consecutive even integers such that four times the sum of the first and second is 8 less than 7 times the fourth 4[x + (x+1)] = 7(x+3) -8 Solving for x will give 9 as the first number, so 9, 10, 11, 12